# Inverse-Free Fast Natural Gradient Descent Method for Deep Learning

### v1

@@ Abstract
Second-order methods can converge much faster than first-order methods by incorporating second-order derivates or statistics, but they are far less prevalent in deep learning due to their computational inefficiency. To handle this, many of the existing solutions focus on reducing the size of the matrix to be inverted. However, it is still needed to perform the inverse operator in each iteration. In this paper, we present a fast natural gradient descent (FNGD) method, which only requires computing the inverse during the first epoch. Firstly, we reformulate the gradient preconditioning formula in the natural gradient descent (NGD) as a weighted sum of per-sample gradients using the Sherman-Morrison-Woodbury formula. Building upon this, to avoid the iterative inverse operation involved in computing coefficients, the weighted coefficients are shared across epochs without affecting the empirical performance.
FNGD approximates the NGD as a fixed-coefficient weighted sum, akin to the average sum in first-order methods. Consequently, the computational complexity of FNGD can approach that of first-order methods. To demonstrate the efficiency of the proposed FNGD, we perform empirical evaluations on image classification and machine translation tasks. For training ResNet-18 on the CIFAR-100 dataset, FNGD can achieve a speedup of 2.05$\times$ compared with KFAC. For training Transformer on Multi30K, FNGD outperforms AdamW by 24 BLEU score while requiring almost the same training time.
@@ I Introduction
First-order methods, such as stochastic gradient descent (SGD) [20], and its adaptive learning rate variants AdaGrad [3], RMSprop [10], and Adam [11], are dominant for the training of deep neural networks (DNNs). Large-scale tasks particularly favor their low complexity property. However, there are some serious issues with first-order methods. Firstly, ill-conditioned objectives may exhibit slow convergence in flat regions and oscillations in steep regions. Secondly, they are sensitive to hyperparameter settings, often necessitating meticulous tuning efforts.
@@
Second-order methods can alleviate these problems. The traditional Newton’s method [17] preconditions the gradient with the inverse of the Hessian matrix. By the preconditioning operator, the gradient can be rotated and rescaled to facilitate escaping from ill-conditioned ’valleys’. However, as DNNs are non-convex, the Hessian matrix does not necessarily be positive semi-definite (PSD). Therefore, practical second-order methods prefer approximations of Hessian that guarantee PSD, particularly the Fisher information matrix (FIM) [1].
@@
In statistical machine learning, the FIM is defined as the covariance of score function. Given a probabilistic model, our objective is to learn its parameters $\mathbf{w}$ by maximizing the log-likelihood. The score function is essentially the gradient of the log-likelihood. As its covariance matrix, i.e. , $\mathbb{E}(\nabla_{\mathbf{w}}\nabla^{\text{T}}_{\mathbf{w}})$, FIM can be equivalent to the negative expected Hessian of the model’s log-likelihood [13]. Consequently, FIM can serve as an alternative to Hessian. However, for a deep learning problem with $N$ parameters, the size of FIM is $N\times N$ and the computational cost for inverting it is $\operatorname{O}(N^{3})$, which is infeasible for DNNs with millions of parameters.
@@
To bridge this gap, a series of algorithms [14, 4, 21] make a block-diagonal approximation to FIM, where blocks correspond to layers. The block approximation cancels out the correlations among layers. Given the low-rank property of FIM, some other methods [19, 21] utilize Sherman-Morrison-Woodbury (SMW) formula [8] to efficiently compute the inverse. However, these methods still can’t avoid performing the inverse operator in each iteration. As a result, the end-to-end training time may approach or even surpass that of SGD.
@@
To address this issue, this work presents a fast natural gradient descent (FNGD) method in which the inverse operator is performed exclusively during the first epoch. Firstly, We find that the gradient preconditioning formula in natural gradient descent (NGD) reformulated by SMW is a matrix-vector multiplication. It has the interpretation of a weighted sum of per-sample gradients. Meanwhile, by re-arranging the computation order, we decrease the preconditioning computational complexity from $\operatorname{O}(N_{l}M^{2}+N_{l}^{2}M+N_{l}^{2})$ to $\operatorname{O}(M^{2}+N_{l}M)$, where $N_{l}$ represents the number of parameters in layer $l$ and $M$ represents the batch size. Moreover, we find that the weighted coefficient vector is solely determined by a correlation matrix that reveals the correlation of samples. This observation inspires us to share them across epochs, and the second-order preconditioning step in NGD is approximated with a fixed-coefficient weighted sum, akin to the average sum in SGD. As a result, the computational complexity of FNGD can approach that of SGD. We provide a computational complexity comparison between FNGD [^1] We ignore the minor computational cost associated with computing coefficients in the first epoch. and conventional second-order methods in Tab. I.
@@
TABLE I: Computational complexity comparison of different optimization algorithms. $D$ is the dimension of a hidden layer, and $M$ is the batch size.
Method Statistics Inverse Precondition
KFAC [14] $\operatorname{O}(2MD^{2})$ $\operatorname{O}(2D^{3})$ $\operatorname{O}(2D^{3})$
Eva [25] $\operatorname{O}(2MD)$ $\operatorname{O}(2D^{2})$
Shampoo [7] $\operatorname{O}(2D^{3})$ $\operatorname{O}(2D^{3})$ $\operatorname{O}(2D^{3})$
FNGD - - $\operatorname{O}(MD^{2})$
@@
We conduct numerical experiments on image classification and machine translation tasks to demonstrate the effectiveness and efficiency of FNGD. In the task of image classification, FNGD can yield comparable convergence and generalization performance as conventional second-order methods, such as KFAC [14], Shampoo [7], and Eva [25]. Moreover, FNGD outperforms them in terms of per-epoch training time. Specifically, compared with KFAC, Shampoo, and Eva, FNGD can achieve up to 2.05$\times$, 1.22$\times$, and 1.44$\times$ time reduction, respectively. For the machine translation task with the Transformer, FNGD outperforms AdamW by 24 BLEU score on Multi30K while requiring almost the same training time. Furthermore, when compared with other second-order methods, FNGD is approximately 2.4$\times$ faster than KFAC and 5.7$\times$ faster than Shampoo.
@@
  • We reformulate the gradient preconditioning formula as a weighted sum of per-sample gradients using the SMW formula. It establishes a connection between NGD and SGD.
  • We share these weighted coefficients across epochs, as they are solely determined by the correlation of samples. As a result, we don’t have to perform the inverse operator involved in computing coefficients in every iteration, only in the initial epoch.
  • The per-sample gradient is efficiently computed by applying AotoGrad to the module output, rather than the module parameters. It can eliminate the needless computations of the average gradient of parameters.
  • The neural networks training experiments illustrate that when compared with KFAC, Shampoo, and Eva, our method can achieve superior computational efficiency while yielding comparative convergence and generalization performance.
@@ II Related Work
Natural gradient methods rely on the FIM to precondition the gradient. To make it practical for DNNs, plenty of works apply a block-diagonal approximation to the FIM. Beyond this basic approximation, additional strategies were proposed to further reduce the computational complexity. There are primarily two approaches: making further approximation or reducing the inverse complexity. In Fig. 2, we outline several methods that make a further approximation on FIM, including KFAC [14], Shampoo [7], Eva [25], and MBF [2]. For the purpose of efficiently computing the inverse, SKFAC [21] proposed to employ the SMW formula to reduce the size of the matrix to be inverted. Moreover, HyLo [16] can further reduce the matrix size by extracting key training samples, enhancing the scalability of KFAC on distributed platforms.
@@
The above Fisher approximation methods focus on computing the statistics or the inverse with low costs. However, they do not analyze the structure of preconditioned gradients. FNGD explores this to cleverly reduce the computational complexity.
@@ III-A Notation
Scalars are denoted by letters, e.g., $a$. Vectors are denoted by boldface lowercase letters, e.g., $\mathbf{a}$. Matrices are denoted by boldface capital letters, e.g., $\mathbf{A}$. Higher-order tensors are denoted by Euler script letters, e.g., $\mathcal{X}$. The symbol $\odot$ signifies the Khatri-Rao product; $\otimes$ signifies the Kronecker product; *** signifies the Hadamard product; and $||\cdot||_{\text{F}}$ signifies the Frobenius norm of a matrix. The operator $\text{vec}(\cdot)$ reshapes a matrix into a vector by stacking rows.
@@ III-C Natural Gradient Method
The pairs $(\mathbf{x}_{m},\mathbf{y}_{m})$ are from the training dataset. Defining a matrix $\mathbf{U}=[\frac{\partial\operatorname{\ell}(\operatorname{f}(\mathbf{x}_{1},\mathbf{w}),\mathbf{y}_{1})}{\partial\mathbf{w}},\frac{\partial\operatorname{\ell}(\operatorname{f}(\mathbf{x}_{2},\mathbf{w}),\mathbf{y}_{2})}{\partial\mathbf{w}},\cdots,\frac{\partial\operatorname{\ell}(\operatorname{f}(\mathbf{x}_{M},\mathbf{w}),\mathbf{y}_{M})}{\partial\mathbf{w}}]\in\mathbb{R}^{N\times M}$, named the Jacobian matrix, the EFM can be represented in terms of $\mathbf{U}$ as follows:
@@ IV Proposed Method
Firstly, based on the well-known SMW formula, we restructure the updating formula. In this way, we can interpret the preconditioned gradient as a weighted sum of per-sample gradients. By re-arranging computation order, we can decrease the preconditioning computational complexity from $\operatorname{O}(N_{l}M^{2}+N_{l}^{2}M+N_{l}^{2})$ to $\operatorname{O}(M^{2}+N_{l}M)$. Furthermore, by sharing these weighted coefficients across epochs, we approximate the preconditioning step in NGD as a fixed-coefficient weighted sum, which closely resembles the average sum in SGD. Consequently, the theoretical complexity of FNGD is comparable to that of SGD. For the implementation, we provide a discussion on how to efficiently compute the per-sample gradient.
@@ IV-B Coefficient-Sharing
In Eq. 9, we represent the preconditioning equation as a matrix-vector multiplication, highlighting that the preconditioned gradient is a weighted sum of per-sample gradients within a mini-batch.
@@
Delving further, we can see that the weighted coefficient vector is solely determined by the matrix $\mathbf{U}_{l}^{\text{T}}\mathbf{U}_{l}$. The matrix $\mathbf{U}_{l}^{\text{T}}\mathbf{U}_{l}$ is, in essence, a Gram matrix, where each entry represents the similarity between the gradients with respect to two training samples. In this way, we figure that $\mathbf{U}_{l}^{\text{T}}\mathbf{U}_{l}$ reveals the correlation of $M$ samples. If there are two distinct samples, the direction of their gradients may be orthogonal in the parameter space. Consequently, the corresponding entry of $\mathbf{U}_{l}^{\text{T}}\mathbf{U}_{l}$ may be close to zero. On the contrary, for two similar samples, the gradients tend to align closely in direction, resulting in a large entry of $\mathbf{U}_{l}^{\text{T}}\mathbf{U}_{l}$.
@@
For each layer, we have individual correlation matrix $\mathbf{U}_{l}^{\text{T}}\mathbf{U}_{l}$. This setting aligns with the concept of “hierarchical feature learning” in deep learning [5]. As training samples pass through the deep network, lower layers tend to capture basic features, while higher layers capture more abstract and complex features. Consequently, the correlation matrix of the higher layer is expected to differ significantly from that of the lower layer, whereas the adjacent layers may exhibit similar correlation matrices. We depict this point in Fig. 3.
@@
Figure 3: The correlation matrix $\mathbf{U}_{l}^{\text{T}}\mathbf{U}_{l}$ for four layers in ResNet-32 [9] on CIFAR-10 with batch size 128. The first two layers are adjacent, as are the last two layers.
The above observation on $\mathbf{U}_{l}^{\text{T}}\mathbf{U}_{l}$ inspires us to apply coefficient-sharing. Neglecting the subtle influence of random data augmentation over epochs, the training samples used in each epoch remain constant. Therefore, the correlation matrix, determined by the intrinsic nature of samples, is considered to be unchanging. We can ignore the influence of gradient magnitude on the correlation matrix, as the variation in magnitude can be canceled out by multiplying $\mathbf{U}_{l}^{\text{T}}\mathbf{g}_{l}$. As a result, we make the hypothesis that the weighted coefficients, which solely depend on the correlation matrix, are constant across epochs. This leads to the technique of coefficient-sharing.
@@
Taking the mini-batch strategy into account, the weighted sum is performed within each mini-batch. Routinely, in order to improve the generalization performance, the training dataset is shuffled before being divided into batches in each epoch. This shuffling operator randomizes the samples within each batch. However, through our empirical experiments, we have found that coefficient-sharing across epochs continues to be effective despite the variability in samples within batches.
@@
In Fig. 4, we demonstrate the comparative results (with or without coefficient-sharing) for ResNet-32 on the Cifar-10 dataset. Remarkably, FNGD, the one with coefficient-sharing, achieves performance that is on par with the baseline NGD. On the other hand, thanks to the technique of coefficient-sharing, there is no need to compute the second-order information for epochs beyond the first epoch. Consequently, this results in a significant reduction in time costs, as depicted in Fig. 4. Specifically, FNGD is shown to be twice as fast as NGD.
@@
To clarify the effectiveness of FNGD, we can interpret the impact of shuffling-induced randomness from a different perspective. The gradient of a sample assigned with a large coefficient exerts a significant impact on the updating process, while that of a sample assigned with a small coefficient has a relatively minor impact. As a result, with coefficients constant, we randomly pick certain samples as key contributors to guide the optimization process. It may increase the model’s robustness to noise in data. However, it’s important to emphasize that we can’t simply set random coefficients without initially computing Eq. 9 during the first epoch. This is because, due to the potential similarity in $\mathbf{U}_{l}^{\text{T}}\mathbf{U}_{l}$ across layers, the coefficients of several layers are likely to be coupled. This level of interdependence among coefficients wouldn’t be achieved through random initialization alone.

### v2

@@ Abstract
Second-order optimization techniques have the potential to achieve faster convergence rates compared to first-order methods through the incorporation of second-order derivatives or statistics. However, their utilization in deep learning is limited due to their computational inefficiency. Various approaches have been proposed to address this issue, primarily centered on minimizing the size of the matrix to be inverted. Nevertheless, the necessity of performing the inverse operation iteratively persists. In this work, we present a fast natural gradient descent (FNGD) method that only requires inversion during the first epoch. Specifically, it is revealed that natural gradient descent (NGD) is essentially a weighted sum of per-sample gradients. Our novel approach further proposes to share these weighted coefficients across epochs without affecting empirical performance. Consequently, FNGD exhibits similarities to the average sum in first-order methods, leading to the computational complexity of FNGD being comparable to that of first-order methods. Extensive experiments on image classification and machine translation tasks demonstrate the efficiency of the proposed FNGD. For training ResNet-18 on CIFAR-100, FNGD can achieve a speedup of 2.07$\times$ compared with KFAC. For training Transformer on Multi30K, FNGD outperforms AdamW by 24 BLEU score while requiring almost the same training time.
@@ I Introduction
First-order methods, such as stochastic gradient descent (SGD) [23], and its adaptive learning rate variants including AdaGrad [5], RMSprop [12], and Adam [13], are dominant for training deep neural networks (DNNs), especially in large-scale tasks. Nonetheless, these methods face significant challenges. One key issue is the slow convergence observed in flat regions or the oscillations observed in steep regions. Additionally, first-order methods are highly sensitive to hyperparameter configurations, often requiring extensive tuning efforts.
@@
Second-order methods offer a solution to these problems [26, 3, 4]. Newton’s method [20], a conventional second-order approach, involves preconditioning gradient with the inverse of Hessian matrix. This preconditioning operation enables gradient to be rotated and rescaled, aiding in the escape from ill-conditioned ’valleys’. However, due to the non-convex nature of DNNs, Hessian matrix may not always be positive semi-definite (PSD). Consequently, in practical applications, second-order methods often rely on approximations of Hessian that ensure PSD properties, such as the Fisher information matrix (FIM) [1].
@@
In the field of statistical machine learning, FIM is defined as the covariance of score function. Given a probabilistic model, the score function represents the gradient of log-likelihood with respect to model parameters $\mathbf{w}$. FIM, serving as the corresponding covariance matrix, i.e. $\mathbb{E}(\nabla_{\mathbf{w}}\nabla^{\text{T}}_{\mathbf{w}})$, has been proved to be equal to the negative expected Hessian of the model’s log-likelihood [16]. Consequently, FIM can serve as an alternative to Hessian. However, in the case of deep learning with $N$ parameters, the FIM matrix size is $N\times N$, and the computational complexity for its inversion is $\operatorname{O}(N^{3})$, rendering it impractical for DNNs with millions of parameters.
@@
To bridge this gap, a series of algorithms [17, 6, 24] have been developed to approximate FIM by employing a block-diagonal structure, with each block corresponding to a specific layer. This block-wise approximation effectively eliminates inter-layer correlations. Furthermore, recognizing the low-rank nature of FIM, advanced methods [22, 24] utilize Sherman-Morrison-Woodbury (SMW) formula [10] to efficiently compute the matrix inverse. However, these techniques still necessitate the inverse operator in each epoch. As a result, the end-to-end training time of NGD may approach or even surpass that of SGD.
@@
To address this issue, this work presents a fast natural gradient descent (FNGD) method in which the inverse operator is performed exclusively during the first epoch. Firstly, we find that the gradient preconditioning formula in natural gradient descent (NGD) can be reformulated by SMW into a matrix-vector multiplication. It has the interpretation of a weighted sum of per-sample gradients. By re-arranging the computation order, we decrease the preconditioning computational complexity from $\operatorname{O}(N_{l}M^{2}+N_{l}^{2}M+N_{l}^{2})$ to $\operatorname{O}(M^{2}+N_{l}M)$, where $N_{l}$ represents the number of parameters in layer $l$ and $M$ represents the batch size. Furthermore, it is analyzed that the weighted coefficient vector is solely determined by a correlation matrix that reveals the correlation of samples. This observation inspires us to share it across epochs, making FNGD akin to the average sum in SGD. As a result, the computational complexity of FNGD can approach that of SGD. We provide a complexity comparison between FNGD [^1] We ignore the minor computational cost associated with computing coefficients in the first epoch. and conventional second-order methods in Tab. I.
@@
TABLE I: Complexity comparison of different optimization algorithms. $D$ is the dimension of a hidden layer, and $M$ is the batch size.
Method Statistics Inverse Precondition Memory
KFAC [17] $\operatorname{O}(2MD^{2})$ $\operatorname{O}(2D^{3})$ $\operatorname{O}(2D^{3})$ $\operatorname{O}(2D^{2})$
Eva [30] $\operatorname{O}(2MD)$ $\operatorname{O}(2D^{2})$ $\operatorname{O}(2D)$
Shampoo [9] $\operatorname{O}(2D^{3})$ $\operatorname{O}(2D^{3})$ $\operatorname{O}(2D^{3})$ $\operatorname{O}(2D^{2})$
FNGD - - $\operatorname{O}(MD)$ $\operatorname{O}(M)$
@@
We conduct numerical experiments on image classification and machine translation tasks to demonstrate the effectiveness and efficiency of FNGD. In the task of image classification, FNGD can yield comparable convergence and generalization performance to conventional second-order methods, such as KFAC [17], Shampoo [9], and Eva [30]. Additionally, FNGD surpasses these methods in terms of computational efficiency. Specifically, in comparison to KFAC, Shampoo, and Eva, FNGD can achieve time reductions of up to 2.65$\times$, 2.03$\times$, and 1.73$\times$, respectively. In the context of machine translation with Transformer, FNGD outperforms AdamW by 24 BLEU score on Multi30K while requiring almost the same training time. Furthermore, when compared with other second-order methods, FNGD is approximately 2.4$\times$ faster than KFAC and 5.7$\times$ faster than Shampoo.
@@
  • We reformulate the gradient preconditioning formula of NGD as a weighted sum of per-sample gradients using SMW formula. It establishes a connection between NGD and SGD.
  • We propose sharing these weighted coefficients across epochs by analyzing the physical meaning of coefficients. This approach eliminates the need to perform inverse operator for coefficients computation in every iteration, except in the initial epoch.
  • An efficient preconditioning strategy is proposed to accelerate the implementation of FNGD, including efficient backward and weighted sum.
  • we apply our approach for image classification and machine translation tasks to demonstrate its time efficiency.
@@ II Related Work
Natural gradient methods rely on FIM to precondition the gradient. To make it practical for DNNs, plenty of works apply a block-diagonal approximation to the FIM. In addition to this fundamental approximation, various techniques have been introduced to further reduce the computational complexity. There are primarily two approaches: making further approximation or reducing the inverse complexity. In Fig. 2, we outline several methods that offer further approximations of FIM, including KFAC [17], Shampoo [9], Eva [30], and MBF [2]. To enhance the efficiency of computing the inverse, SKFAC [24] proposed to employ the SMW formula to decrease the size of the matrix to be inverted. Furthermore, HyLo [19] has the capability to decrease the matrix size by selecting important training samples, thereby improving the scalability of KFAC on distributed platforms.
@@
The Fisher approximation techniques discussed above primarily focus on computing the statistics or the inverse with low costs. However, these methods do not delve into analyzing the structure of preconditioned gradients. FNGD, on the other hand, strategically addresses this aspect to reduce the computational complexity effectively.
@@ III-A Notation
Scalars are denoted by letters, e.g., $a$. Vectors are denoted by boldface lowercase letters, e.g., $\mathbf{a}$. Matrices are denoted by boldface capital letters, e.g., $\mathbf{A}$. Higher-order tensors are denoted by Euler script letters, e.g., $\mathcal{X}$. The symbol $\odot$ signifies the Khatri-Rao product; $\otimes$ signifies the Kronecker product; * signifies the Hadamard product; and $||\cdot||_{\text{F}}$ signifies the Frobenius norm of a matrix. The operator $\text{vec}(\cdot)$ reshapes a matrix into a vector by stacking rows.
@@ III-C Natural Gradient Method
The pairs $(\mathbf{x}_{m},\mathbf{y}_{m})$ are from the training dataset. Defining a matrix $\mathbf{U}=[\frac{\partial\operatorname{\ell}(\operatorname{f}(\mathbf{x}_{1},\mathbf{w}),\mathbf{y}_{1})}{\partial\mathbf{w}},\frac{\partial\operatorname{\ell}(\operatorname{f}(\mathbf{x}_{2},\mathbf{w}),\mathbf{y}_{2})}{\partial\mathbf{w}},\cdots,\frac{\partial\operatorname{\ell}(\operatorname{f}(\mathbf{x}_{M},\mathbf{w}),\mathbf{y}_{M})}{\partial\mathbf{w}}]\in\mathbb{R}^{N\times M}$, named the Jacobian matrix of the loss, the EFM can be represented in terms of $\mathbf{U}$ as follows:
@@ IV Proposed Method
Firstly, based on the well-known SMW formula, we propose to restructure the updating formula in Eq. 4. This adjustment allows for the interpretation of the preconditioned gradient as a weighted sum of per-sample gradients. By re-arranging computation order, we can decrease the preconditioning computational complexity from $\operatorname{O}(N_{l}M^{2}+N_{l}^{2}M+N_{l}^{2})$ to $\operatorname{O}(M^{2}+N_{l}M)$. Furthermore, by sharing these weighted coefficients across epochs, we approximate the preconditioning step in NGD as a fixed-coefficient weighted sum, which closely resembles the average sum in SGD. Consequently, the theoretical complexity of FNGD is comparable to that of SGD. For the implementation, we provide a discussion on how to compute the per-sample gradient and perform the weighted sum efficiently. The framework of FNGD is depicted in Algorithm 1.
Algorithm 1 Framework of FNGD
0:  Learning rate: $\eta$; Damping value: $\lambda$; Batch size: $M$; number of batches: $B$. // Derive coefficients during the first epoch
1:  for $i=1$ to $B$ do
2:     Sample mini-batch of size $M$;
3:     Perform forward-backward pass to compute $\mathbf{U}_{i}$;
4:     Compute coefficients $\mathbf{v}_{i}$ based on Eq. 9;
5:     Perform parameter update: $\mathbf{w}\leftarrow\mathbf{w}-\eta\mathbf{U_{i}}\mathbf{v}_{i}$;
6:  end for
7:  Compute average coefficient: $\tilde{\mathbf{v}}=\frac{1}{B}\sum_{i=1}^{B}\mathbf{v}_{i}$;
8:  while not converage do
9:     for $i=1$ to $B$ do
10:        Sample mini-batch of size $M$;
11:        Perform forward-backward pass to derive gradients of layers’ output $\mathcal{A}$;
12:        Apply weighted coefficients on layers’ inputs $\mathcal{B}$;
13:        $\tilde{\mathbf{g}}\leftarrow$ torch.nn.grad.layer_weight($\mathcal{A}$, $\mathcal{B}$);
14:        Perform parameter update: $\mathbf{w}\leftarrow\mathbf{w}-\eta\tilde{\mathbf{g}}$;
15:     end for
16:  end while
@@ IV-B Coefficient-Sharing
In Eq. 9, we represent the preconditioning equation as a matrix-vector multiplication, highlighting that the preconditioned gradient is a weighted sum of per-sample gradients. Building upon this reformulation, we propose reasonably sharing the coefficient across iterations to avoid the computationally intensive inverse operation.
@@
Firstly, we can see that the weighted coefficient vector is solely determined by the matrix $\mathbf{U}_{l}^{\text{T}}\mathbf{U}_{l}$. The matrix $\mathbf{U}_{l}^{\text{T}}\mathbf{U}_{l}$ is, in essence, a Gram matrix, where each entry reveals the similarity between two training samples. Specifically, when two samples are distinct, their gradient directions might be orthogonal within the parameter space, leading to a near-zero entry in $\mathbf{U}_{l}^{\text{T}}\mathbf{U}_{l}$. In contrast, gradients for similar samples are likely to closely align in direction, yielding a significant entry in $\mathbf{U}_{l}^{\text{T}}\mathbf{U}_{l}$.
@@
For each layer, we have an individual correlation matrix $\mathbf{U}_{l}^{\text{T}}\mathbf{U}_{l}$. This setting aligns with the concept of “hierarchical feature learning” in deep learning [7]. As training samples pass through the deep network, lower layers tend to capture basic features, while higher layers capture more abstract and complex features. Consequently, the correlation matrix of the higher layer is expected to differ significantly from that of the lower layer, whereas the adjacent layers may exhibit similar correlation matrices. We depict this point in Fig. 3.
@@
Figure 3: The normalized correlation matrix $\mathbf{U}_{l}^{\text{T}}\mathbf{U}_{l}$ for four layers in ResNet-32 [11] on CIFAR-10 with batch size 128. The first two layers are adjacent, as are the last two layers.
From an optimization perspective, the gradient weighted sum technique is an analogy to the weighted loss concept [28, 14], which is widely employed in machine learning to direct the optimization’s attention. With fixed weighting coefficients, samples are assigned varying degrees of importance. However, a unique aspect of FNGD lies in its layer-specific optimization approach, where each layer is assigned distinct weighting coefficients. This characteristic might contribute to layer-specific feature learning.
Rather than randomly set coefficients, FNGD derives the coefficients of each layer by initially computing Eq. 9 during the first epoch. By doing so, the coefficients of adjacent layers could be coupled. This level of interdependence among coefficients wouldn’t be achieved through random initialization alone.
@@
In practical applications, considering the mini-batch strategy, the weighted sum is performed within each mini-batch. Routinely, in order to enhance the generalization performance, the training dataset is shuffled before being divided into batches in each epoch. This shuffling operator randomizes the samples within each batch. However, our empirical experiment has found that coefficient-sharing across epochs remains effective despite the variability in samples within each batch.
@@
In Fig. 4, we demonstrate the comparative results (with or without coefficient-sharing) for ResNet-32 on the Cifar-10 dataset. Remarkably, FNGD, the one with coefficient-sharing, achieves performance that is on par with the baseline NGD. On the other hand, thanks to the technique of coefficient-sharing, there is no need to compute the second-order information for epochs beyond the first epoch. Consequently, this results in a significant reduction in time costs. Specifically, FNGD is shown to be twice as fast as NGD.
@@
To clarify the effectiveness of FNGD, we can interpret the impact of shuffling-induced randomness from a different perspective. A sample with a large coefficient significantly influences the update process, while one with a small coefficient has a minor impact. As a result, with coefficients constant, FNGD randomly picks certain samples as key contributors to guide the optimization process. It may increase the model’s robustness to noise.